### v2

# DNA-DetectLLM: Unveiling AI-Generated Text via a DNA-Inspired Mutation-Repair Paradigm

Abstract
The rapid advancement of large language models (LLMs) has blurred the line between AI-generated and human-written text. This progress brings societal risks such as misinformation, authorship ambiguity, and intellectual property concerns, highlighting the urgent need for reliable AI-generated text detection methods. However, recent advances in generative language modeling have resulted in significant overlap between the feature distributions of human-written and AI-generated text, blurring classification boundaries and making accurate detection increasingly challenging. To address the above challenges, we propose a DNA-inspired perspective, leveraging a repair-based process to directly and interpretably capture the intrinsic differences between human-written and AI-generated text. Building on this perspective, we introduce DNA-DetectLLM, a zero-shot detection method for distinguishing AI-generated and human-written text. The method constructs an ideal AI-generated sequence for each input, iteratively repairs non-optimal tokens, and quantifies the cumulative repair effort as an interpretable detection signal. Empirical evaluations demonstrate that our method achieves state-of-the-art detection performance and exhibits strong robustness against various adversarial attacks and input lengths. Specifically, DNA-DetectLLM achieves relative improvements of 5.55% in AUROC and 2.08% in F1 score across multiple public benchmark datasets. Code and data are available at https://github.com/Xiaoweizhu57/DNA-DetectLLM.
1 Introduction
The rapid advancement of large language models (LLMs) has created increasingly human-like textual content, substantially narrowing the distinguishable gap between AI-generated and human-written text. While these improvements have catalyzed significant technological breakthroughs, they simultaneously pose critical societal challenges, including misinformation dissemination, authorship ambiguity, and threats to intellectual property rights ahmed2021detectingfakenewsusing ; Adelani2019GeneratingSF ; 9121286 . Consequently, there is an urgent and growing need for effective and reliable methods to accurately detect AI-generated text.
While significant research efforts have been dedicated to AI-generated text detection, existing methodologies typically adopt either training-based or training-free methods. Training-based methods solaiman2019release ; NEURIPS2023_30e15e59 ; NEURIPS2024_a117a3cd ; NEURIPS2024_1d35af80 ; NEURIPS2024_bc808cf2 depend upon large volumes of annotated data, limiting their scalability and generalization to new domains. In contrast, training-free approaches pmlr-v202-mitchell23a ; bao2024fastdetectgpt ; hans2024spotting ; DBLP:journals/corr/abs-2410-06072 leverage intrinsic statistical differences to distinguish human-written and AI-generated texts. Both paradigms fundamentally operate by attempting to identify distinct, separable boundaries within the feature space. However, recent advancements in generative language modeling have produced outputs increasingly indistinguishable from human-authored content, causing these classification boundaries to become progressively blurred. Empirical studies dugan-etal-2024-raid ; sadasivan2025aigeneratedtextreliablydetected have highlighted substantial overlap regions in the feature distributions of human-written and AI-generated texts, significantly undermining detection accuracy in practical scenarios. Therefore, one capable of more precisely and intrinsically capturing differences between the generative processes of AI and human writing is urgently needed.
In molecular biology, DNA’s double-helix structure ensures stable transmission of genetic information, yet mutations during replication introduce variations that can lead to individual differences or even diseases such as cancer. In a similar vein, an ideal AI-generated text sequence can be seen as a “template strand”, representing the most probable token choices at each position. Human-written texts, by contrast, resemble mutated strands, where token selections deviate from the optimal probabilities, creating measurable differences. Inspired by this biological mechanism, we propose a new perspective for AI-generated text detection: by analogizing to DNA base-repair processes, we iteratively “correct” non-optimal tokens in a text and measure the difficulty of restoring it to the ideal AI-generated form. This repair-based approach captures the intrinsic divergence between AI-generated and human-written texts in a direct and interpretable manner.
Building on this intuition, we propose DNA-DetectLLM, a novel method for zero-shot detection of AI-generated texts. For each input sequence, we first construct its corresponding ideal AI sequence—that is, the sequence formed by greedily selecting the most probable token at each position under a reference language model. We then perform a token-by-token repair process on the input sequence, progressively modifying tokens toward their optimal choices until the sequence fully aligns with the ideal AI sequence. To quantify the difficulty of this repair process, we introduce a repair score that captures the cumulative effort required to complete the transformation. Finally, by comparing the repair score against a calibrated threshold, DNA-DetectLLM robustly distinguishes AI-generated texts from human-written ones, leveraging the fundamental differences in their deviation patterns from ideal generation.
Figure 1: Illustration of the Mutation-Repair Paradigm. The input sequence can be analogized to a replicating strand, while the ideal AI-generated sequence corresponds to the template strand.
DNA-DetectLLM consistently achieves state-of-the-art performance across multiple datasets and LLMs. In particular, it obtains relative improvements of 5.55% in AUROC and 2.08% in F1 score on three public benchmark datasets. Additionally, the method exhibits notable robustness against various adversarial attacks and across different input lengths. Efficiency experiments further indicate rapid detection capability, processing each sample in under 0.8s.
Our contributions are summarized as follows:
  • Inspired by the mutation and repair mechanisms of nucleotide bases in DNA replication, we introduce the mutation-repair paradigm into AI-generated text detection.
  • We propose DNA-DetectLLM, a novel zero-shot method for detecting AI-generated text that incrementally repairs mutated tokens within the input sequence until it perfectly aligns with the ideal AI-generated sequence, subsequently quantifying the repair difficulty as a metric for text detection.
  • Extensive evaluations validate that DNA-DetectLLM offers a reliable, efficient, and broadly generalizable solution for AI-generated text detection, with consistent gains across various detection settings.
2 Related Works
Detecting AI-generated text is essential for enhancing public trust and preventing misuse, driving growing interest from both academia and industry. Beyond watermarking techniques DBLP:journals/corr/abs-2301-10226 , which embed identifiable markers during generation, current post hoc detection methods are broadly categorized into training-based and training-free methods.
Training-based Methods.
Such approaches typically involve training classification models to distinguish between AI-generated and human-written texts. Specifically, early efforts by OpenAI solaiman2019release employed RoBERTa-based models for training text classifiers. Subsequently, RADAR NEURIPS2023_30e15e59 introduced adversarial learning to enhance the robustness against paraphrased texts. DeTeCtive NEURIPS2024_a117a3cd utilized multi-level contrastive learning to map texts generated by different LLMs into corresponding feature spaces, classifying them based on similarity metrics. DPIC NEURIPS2024_1d35af80 extracted deep textual features by reconstructing prompts and regenerating texts. Biscope NEURIPS2024_bc808cf2 proposed employing a bidirectional cross-entropy loss to extract statistical features for binary classifier training. R-Detect song2025deep employs a nonparametric kernel relative test to detect AI-generated text, thereby reducing the false positive rate compared to two-sample tests. However, existing research chakraborty2023possibilitiesaigeneratedtextdetection ; uchendu-etal-2020-authorship indicates that training-based methods consistently overfit to in-distribution features, resulting in poor generalization to out-of-distribution (OOD) texts. Consequently, researchers have increasingly focused on developing more universally applicable training-free methods.
Training-free Methods.
These training-free methods emphasize exploiting probabilistic characteristics of texts, constructing statistical scores based on specific hypotheses, and making decisions according to the comparison of scores against thresholds. For example, LogRank gehrmann-etal-2019-gltr , Likelihood hashimoto-etal-2019-unifying , and Entropy ippolito-etal-2020-automatic calculate the average probability ranking, likelihood probabilities, and entropy values to measure the uncertainty of AI-generated texts. DetectGPT pmlr-v202-mitchell23a pioneered a paradigm that uses perturbations to generate numerous contrast samples to evaluate the overall distribution. Although methods such as DetectLLM-NPR su2023detectllmleveraginglogrank and DNA-GPT DBLP:conf/iclr/Yang0WPWC24 have further developed this paradigm, their efficiency limitations prevent real-time or large-scale detection implementations. Fast-DetectGPT bao2024fastdetectgpt has since updated sampling techniques to compute conditional probability curvature, significantly improving detection efficiency and broadening potential applications. Binoculars hans2024spotting achieved state-of-the-art classification performance by calculating cross-perplexity from dual-model perspectives. Lastde++ DBLP:journals/corr/abs-2410-06072 proposed focusing on local textual features by calculating Diversity Entropy to optimize classification performance.
3 DNA-DetectLLM
3.1 Preliminary
This study primarily involves two statistical metrics: log-perplexity, which quantifies the average token-level negative log-likelihood under a single model, and cross-perplexity, which captures the average per-token cross-entropy between the probability distributions of two models:
$\log\mathrm{PPL}_{M_{1}}(s)$ $=-\frac{1}{L}\sum_{i=1}^{L}\log P_{M_{1}}(x_{i}|x_{<i}),\quad$ (1)
$\log\mathrm{X\text{-}PPL}_{M_{1},M_{2}}(s)$ $=-\frac{1}{L}\sum_{i=1}^{L}P_{M_{1}}(x_{i}|x_{<i})\log P_{M_{2}}(x_{i}|x_{<i}),$
where $s$ is the input sequence of length $L$, $x_{i}$ denotes the $i$-th token, and $P_{M}(x_{i}|x_{<i})$ is the conditional probability of $x_{i}$ given its preceding tokens under reference model $M_{1}$ or observer model $M_{2}$. Furthermore, their ratio $\sigma(s)$ has been empirically demonstrated to serve as an effective score for distinguishing AI-generated text hans2024spotting . To quantify the effect of local token-level modifications on these metrics, this work introduces the conditional log-perplexity and conditional score:
$\log\mathrm{PPL}_{M_{1}}(\tilde{s}|s)=-\frac{1}{L}\sum_{i=1}^{L}\log P_{M_{1}}(\tilde{x}_{i}|x_{<i}),\quad\sigma(\tilde{s}|s)=\frac{\log\mathrm{PPL}_{M_{1}}(\tilde{s}|s)}{\log\mathrm{X\text{-}PPL}_{M_{1},M_{2}}(s)},$ (2)
where $\tilde{s}$ denotes the sequence obtained by modifying tokens in the input sequence $s$.
Figure 2: Overview of DNA-DetectLLM.
3.2 Overview of DNA-DetectLLM
The entire workflow of DNA-DetectLLM can be summarized in 3 key steps, shown in Figure 2.
Step 1: Obtaining the ideal AI-generated Sequence. We construct the ideal AI-generated sequence for a given input by greedily selecting the most probable token at each position.
Step 2: Mutation Repair Mechanism. We perform iterative token-level modifications on the input sequence until it fully aligns with the ideal AI-generated sequence.
Step 3: Repair Score-Based Detection. We introduce a repair score to quantify the difficulty of the repair, which is compared against a calibrated threshold to determine the detection result.
3.3 Obtaining the ideal AI-generated Sequence
We propose the concept of an ideal AI-generated sequence $\hat{s}$, analogous to the error-free template strand in DNA replication, where each token is selected by maximizing the conditional probability at its position:
$\hat{s}=\{\hat{x}_{1},\hat{x}_{2},\ldots,\hat{x}_{L}\},\quad\text{where }\hat{x}_{i}=\arg\max_{\tilde{x}\in\mathcal{V}}P_{M_{1}}(\tilde{x}|x_{<i}),$ (3)
with $\mathcal{V}$ denoting the vocabulary and $x_{<i}=\{x_{1},\ldots,x_{i-1}\}$ representing the preceding i1i{-}1 tokens of the input sequence $s$.
3.4 Mutation Repair Mechanism
We treat non-max probability tokens as mutated tokens and max-probability tokens as ideal tokens. Analogous to the mutation-repair paradigm in DNA, we propose a mutation repair mechanism aiming to uncover fundamental differences in deviation patterns between AI-generated and human-written texts. Under this mechanism, mutated tokens in the input sequence are iteratively repaired with their ideal tokens step by step, until the input fully aligns with the ideal form:
$x_{i}\in s=\{x_{1},x_{2},\ldots,x_{L}\}\rightarrow\hat{x}_{i}=\arg\max_{\tilde{x}\in\mathcal{V}}P_{M_{1}}(\tilde{x}|x_{<i}),\quad\text{if }x_{i}\neq\hat{x}_{i},$ (4)
3.5 Repair Score-Based Detection
We introduce the repair score $R(s)$ to quantify the difficulty of the repair process, defined as the average conditional score accumulated throughout the repair trajectory:
$R(s)=\frac{1}{T+1}\sum_{t=0}^{T}\sigma(s_{t}|s)=\frac{\sum_{t=0}^{T}\log\mathrm{PPL}_{M_{1}}(s_{t}|s)}{(T+1)\log\mathrm{X\text{-}PPL}_{M_{1},M_{2}}(s)},$ (5)
where $s_{t}$ is the sequence after $t$ repair steps, and $T$ is the total number of mutated tokens to be corrected.
Human-written texts typically exhibit more substantial mutations, resulting in greater repair difficulty. In contrast, AI-generated texts are generally easier to repair. Accordingly, the detection result for the input sequence is determined as:
$\mathcal{D}(s)=\begin{cases}\text{ Human-written Text, }&R(s)>\tau\\ \text{ AI-generated Text, }&R(s)\leq\tau.\end{cases}$ (6)
Figure 3: Variation of repair scores Across Different Repair Strategies.
3.6 Sensitivity to Repair Order and Score Simplification
Figure 3 shows that different repair orders yield varying repair scores for the same input sequence, due to the unequal impact of each mutated token on the conditional score. Mutated tokens can be broadly categorized into high- and low-probability types. Repairing low-probability tokens typically causes larger shifts in the conditional score, while high-probability tokens lead to smaller changes. To systematically analyze the influence of repair order, we identify four types of principal repair strategies as follows:
  • High-to-low: Repairing high-probability tokens first, followed by low-probability ones, results in a convex “repair curve” with a higher repair score.
  • Low-to-high: Repairing low-probability tokens before high-probability ones yields a concave “repair curve” with a lower repair score.
  • Sequential Repair: Tokens are repaired in their original order of appearance in the input sequence, regardless of their probability values.
  • Random Repair: Tokens are repaired in a randomly chosen order. Averaging the repair scores across multiple random repairs leads to a more stable estimate.
These findings highlight the sensitivity of the repair score to the chosen repair order. Performing multiple random repairs effectively mitigates biases, resulting in a repair score close to the midpoint between the initial and final scores. Consequently, we further derive that the average repair score converges as the number of random repairs $N$ approaches infinity. The derivation is as follows:
[…]
$\sigma(s_{t}|s)=\sigma(s)-\sum_{i=1}^{t}\delta_{i}\quad\text{for }t=1,2,\ldots,T,\quad\text{where }\sigma(s)=\sigma(s_{0}|s).$ (7)
For a specific permutation $\phi\in[1,N]$, the corresponding repair score $R^{\phi}(s)$} is:
${R^{\phi}(s)}=\frac{1}{T+1}\sum_{t=0}^{T}\left(\sigma(s)-\sum_{i=1}^{t}\delta^{\phi}_{i}\right)=\sigma(s)-\frac{1}{T+1}\sum_{i=1}^{T}\delta^{\phi}_{i}\cdot(T-i+1).$ (8)
Since each token is equally likely to be repaired at random, the expected value of $\delta^{\phi}_{i}$ appearing in the $i$-th position across all permutations is $\frac{1}{T}(\sigma(s)-\sigma(\hat{s}|s)$. We further derive the expected value of the repair score as follows:
$\mathbb{E}[{R(s)}]=\sigma(s)-\frac{1}{T+1}\sum_{i=1}^{T}\mathbb{E}\left[\delta^{\phi}_{i}\right]\cdot(T-i+1)=\sigma(s)-\frac{1}{T+1}\cdot\frac{\sigma(s)-\sigma(\hat{s}|s)}{T}\cdot\frac{T(T+1)}{2}.$ (9)
Therefore, as the number of random permutations $N$ approaches infinity, the average repair score converges as follows:
$\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{n=1}^{N}R^{(n)}(s)=\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{n=1}^{N}\left(\frac{1}{T+1}\sum_{t=0}^{T}\sigma^{(n)}(s_{t}|s)\right)=\frac{1}{2}(\sigma(s)+\sigma(\hat{s}|s)).$ (10)
We thus simplify the repair score to $R(s)=\frac{1}{2}(\sigma(s)+\sigma(\hat{s}|s))$, improving detection performance while avoiding intermediate score computations during repair.
4 Experiments
4.1 Experimental Setup
Datasets.
To evaluate performance across diverse domains, we collect 4,800 human-written texts from three representative tasks: news article writing (XSum narayan-etal-2018-dont ), story generation (WritingPrompts fan-etal-2018-hierarchical ), and academic writing (Arxiv paul2021arxiv ). For each text, we construct task-specific prompts (see Appendix C) and generate corresponding AI outputs using three advanced LLMs: GPT-4 Turbo, Gemini-2.0 Flash, and Claude-3.7 Sonnet. We further sample 2,000 balanced examples from each of three high-quality detection benchmarks—M4 wang-etal-2024-m4 , DetectRL wu2024detectrl , and RealDet zhu2025reliablyboundingfalsepositives —to ensure fair and comprehensive evaluation across real-world scenarios.
Metrics.
We adopt the area under the receiver operating characteristic curve (AUROC https://doi.org/10.1111/j.1466-8238.2011.00683.x ) and F1 score to evaluate detection performance, where higher values indicate better separability between human-written and AI-generated texts.
Baselines.
We compare DNA-DetectLLM with existing training-based and training-free methods. For training-based methods, we include OpenAI-D solaiman2019release , Biscope NEURIPS2024_bc808cf2 and R-Detect song2025deep . For training-free methods, we consider classic zero-shot detectors including Likelihood hashimoto-etal-2019-unifying , LogRank gehrmann-etal-2019-gltr , and Entropy ippolito-etal-2020-automatic , along with several recent SOTA approaches such as DetectGPT pmlr-v202-mitchell23a , Fast-DetectGPT bao2024fastdetectgpt , Binoculars hans2024spotting , and Lastde++ DBLP:journals/corr/abs-2410-06072 . More baseline comparisons are provided in Appendix E.
Implementation details.
In real-world detection scenarios, the source and distribution of textual data are often unknown, constituting an out-of-distribution (OOD) detection problem. To ensure fairness for training-based methods, we exclusively train on the HC3 dataset guo2023closechatgpthumanexperts , which is entirely disjoint from the test sets. For training-free methods, the choice of LLM used for scoring can introduce significant performance variation bao2025glimpse . To eliminate this factor, we standardize the reference (or scoring) model across all methods by employing Falcon-7B-Instruct refinedweb to compute token generation probabilities. Moreover, Fast-DetectGPT, Binoculars, Lastde++, and DNA-DetectLLM utilize Falcon-7B refinedweb as the observer (or sampling) model, while DetectGPT uses T5-3B 2020t5 . During testing, the maximum input token length is capped at 1024. More details are in Appendix D.
Table 1: AUROC (%) of detectors on human-written vs. AI-generated text across datasets and LLMs.
Detectors XSum WritingPrompt Arxiv Avg.
GPT-4
Turbo
Gemini-2.0
Flash
Claude-3.7
Sonnet
GPT-4
Turbo
Gemini-2.0
Flash
Claude-3.7
Sonnet
GPT-4
Turbo
Gemini-2.0
Flash
Claude-3.7
Sonnet
Training-based Methods
OpenAI-D 60.51 68.93 62.96 50.94 59.47 57.28 49.63 51.40 68.57 58.85
Biscope 75.08 95.63 94.09 80.08 98.71 98.05 82.53 99.74 96.61 91.17
R-Detect 63.56 45.63 51.13 73.58 71.07 75.74 56.47 57.24 53.55 60.89
Training-free Methods
Entropy 72.26 54.85 74.90 87.85 90.36 91.17 45.60 79.96 80.42 75.26
Likelihood 70.03 69.50 70.39 80.82 95.52 85.85 57.87 93.60 86.24 78.87
LogRank 69.61 69.26 69.81 78.90 94.53 84.13 58.17 94.15 85.80 78.26
DetectGPT 61.50 68.09 61.36 72.80 89.33 78.17 56.10 92.18 90.27 74.42
Fast-DetectGPT 98.33 92.54 94.30 97.79 98.58 94.14 92.12 99.80 98.21 96.20
Binoculars 98.06 95.56 96.83 97.73 99.53 97.29 93.69 99.87 97.99 97.39
Lastde++ 97.25 90.87 92.77 95.12 98.19 92.36 90.43 99.54 97.57 94.90
DNA-DetectLLM 99.31 96.65 98.45 98.86 99.72 98.51 95.00 99.88 98.35 98.30
DNA-DetectLLM with Other Repair Orders
Low-to-high 98.94 94.98 97.31 98.29 99.41 97.77 92.48 99.75 97.91 97.43
High-to-low 99.14 97.22 98.40 98.63 99.81 98.37 94.67 99.77 97.44 98.16
Sequential Repair 98.90 97.82 98.37 96.80 99.77 98.55 95.78 99.91 98.21 98.23
Table 2: Detection performance (AUROC and F1 score) on public benchmark datasets.
Detectors M4
DetectRL Multi-LLM
DetectRL Multi-Domain
RealDet Avg.
AUROC $\text{F}_{1}$ AUROC $\text{F}_{1}$ AUROC $\text{F}_{1}$ AUROC $\text{F}_{1}$ AUROC $\text{F}_{1}$
OpenAI-D 77.51 71.18 78.15 71.90 74.60 70.03 84.75 77.47 78.75 72.65
Biscope 79.74 73.08 79.97 73.20 76.52 71.64 92.88 86.90 82.28 76.21
R-Detect 61.91 67.14 67.40 66.56 79.19 73.38 65.93 67.72 68.61 68.70
Entropy 83.72 79.10 64.30 71.92 47.82 69.24 75.42 74.72 67.82 73.75
Likelihood 85.77 78.38 66.82 66.71 48.96 66.69 85.35 79.75 71.73 72.88
LogRank 87.50 80.70 67.30 66.71 50.55 66.69 86.28 80.69 72.91 73.70
DetectGPT 73.13 70.11 49.57 66.67 34.67 66.67 78.69 73.80 59.02 69.31
Fast-DetectGPT 89.77 84.12 82.26 75.93 74.98 68.91 93.25 90.00 85.07 79.74
Binoculars 90.00 87.40 83.21 82.87 77.45 80.20 93.64 90.51 86.08 85.25
Lastde++ 91.43 84.97 75.36 69.24 67.30 66.67 93.90 89.41 82.00 77.57
DNA-DetectLLM 91.74 87.72 88.97 84.85 88.23 84.94 94.48 90.58 90.86 87.02
4.2 Main Results
Table 1 compares the detection performance of DNA-DetectLLM against other baselines across different writing tasks and various generation models. DNA-DetectLLM consistently achieves state-of-the-art performance under all settings, with an average AUROC of 98.30%, representing a relative improvement of 0.93%. Specifically, it yields relative gains of 1.36%, 0.87%, and 0.58% on the XSum, WritingPrompts, and Arxiv datasets, respectively, demonstrating strong cross-domain generalization. This strong generalization can be attributed to DNA-DetectLLM’s ability to dynamically capture generation discrepancies between domain-specific text and its ideal AI-generated counterpart through the mutation-repair mechanism, enabling robust identification of human-written versus AI-generated text across diverse domains.
Table 2 evaluates the real-world detection performance of all methods on three high-quality public benchmarks. DNA-DetectLLM demonstrates superior reliability, with average AUROC and F1 score improvements of 5.55% and 2.08%, respectively. Notably, it achieves significant AUROC gains on the challenging DetectRL settings—6.92% on Multi-LLM and 13.92% on Multi-Domain. This improvement can be attributed to the inherent difficulty of DetectRL, where both positive and negative samples may include mixtures of texts with the same label due to the dataset’s construction. Such complexity hinders traditional training-free methods that rely on fixed statistical scores. In contrast, DNA-DetectLLM accurately computes repair scores for intricately constructed input texts by aligning them with their respective ideal AI-generated sequences. This flexible repair-based scoring allows for more accurate detection under distributional overlap and ambiguous cases, underscoring the practical utility of DNA-DetectLLM in complex detection scenarios.
4.3 Robustness
4.3.1 Robustness against Various Attacks
Figure 4: AUROC curves of DNA-DetectLLM and baselines under paraphrasing and editing attacks.
Figure 5: Detection performance (AUROC) on input texts truncated to the target number of tokens.
Figure 4 illustrates the AUROC curves of DNA-DetectLLM and other baselines against various attacks (see Appendix F for details). We conducted experiments on texts generated by GPT-4 Turbo, Gemini-2.0 Flash, and Claude-3.7 Sonnet, incorporating two distinct attacks: token-level edits and paraphrasing. Editing attacks involved random insertion, deletion, or substitution of tokens at rates of 1%. The paraphrasing attacks employed DIPPER NEURIPS2023_575c4500 to rephrase AI-generated texts. To maintain clean labels, attacks were exclusively applied to AI-generated texts.
The results demonstrate that DNA-DetectLLM exhibits strong robustness against a variety of adversarial attacks. For instance, on GPT-4 Turbo-generated text, our method achieves relative AUROC improvements of 6.65%, 3.17%, 6.62%, and 0.81% under insertion, deletion, substitution, and paraphrasing attacks, respectively. Notably, the improvement is particularly pronounced under low false positive rate (FPR) conditions. We attribute this robustness to the observation that although token-level edits are limited in scope, they can substantially alter the generation probability distribution of the input sequence while having minimal impact on its ideal AI-generated sequence. As a result, DNA-DetectLLM is still able to compute accurate repair scores for reliable detection. Moreover, even under paraphrasing attacks using Dipper, the method effectively captures intrinsic deviations from the ideal sequence and maintains a high AUROC of 97.23%. These findings highlight DNA-DetectLLM’s capacity to detect adversarially manipulated AI-generated text, even when such attacks are designed to evade detection.
4.3.2 Robustness on Different Lengths
Prior research bao2024fastdetectgpt ; tian2024multiscale indicates that token length significantly affects detection performance, with shorter texts proving more challenging to detect. We investigate the impact by truncating the input texts to various target tokens. Figure 5 presents the detection performance across varying lengths for five methods: DNA-DetectLLM, Binoculars, Fast-DetectGPT, Lastde++, and Biscope. Results show that DNA-DetectLLM consistently outperforms all baselines across varying lengths. On GPT-4 Turbo-generated text, it achieves an average AUROC improvement of 2.46%. While all methods benefit from longer inputs, DNA-DetectLLM exhibits a greater advantage on shorter texts. At a token length of 40, it surpasses the second-best method by 3.38%, 1.80%, and 3.27%, respectively. These findings suggest that our method enables detection at shorter lengths by extracting more discriminative features from limited textual input.
Figure 6: Comparison of DNA-DetectLLM’s performance when using different LLM pairs.
Figure 7: Comparison of time costs for processing a single sample for each method.
4.4 Ablation Studies
We further evaluated the importance of repair order and different base LLMs through two types of ablation experiments. More detailed ablation studies are available in Appendix G.
Repair Score-based Detection under Various Repair Orders. Table 1 compares the detection performance of DNA-DetectLLM under various repair orders. When the repair order is changed to High-to-low, Low-to-high, or Sequential Repair, a slight performance drop is observed. Although these strategies still outperform other baselines, they require recalculating the conditional score after each mutated token repair, resulting in significantly increased computational cost. In contrast, the simplified repair score ($R(s)=\frac{1}{2}(\sigma(s)+\sigma(\hat{s}|s))$) maintains strong performance while improving efficiency by an order of magnitude, highlighting its necessity in practical deployment.
DNA-DetectLLM’s Performance with Different $M_{1}$ and $M_{2}$. Figure 7 evaluates four different LLM combinations: “Falcon-7B-Instruct + Falcon-7B”, “Llama-3-8B-Instruct + Llama-3-8B”, “Mistral-7B-Instruct + Mistral-7B”, and “Llama-2-7B + Llama-7B”. Results demonstrate that any of these combinations significantly outperform existing baselines, with an average performance improvement of 15.28%. Interestingly, the combination “Llama-2-7B + Llama-7B” slightly exceeds the default combination “Falcon-7B-Instruct + Falcon-7B” used in our main experiments, achieving AUROC of 92.4% and 90.7%. These findings highlight the inherent effectiveness of DNA-DetectLLM, suggesting its robust detection performance is not reliant on any specific LLM combination, with potential for further enhancement through better LLM pairings.
4.5 Efficiency Analysis
Efficiency is critical for AI-generated text detection, as slow detection speeds hinder large-scale or real-time monitoring in practical scenarios. Figure 7 illustrates the average processing time per sample for each method. To eliminate the confounding factor of text length, we randomly sampled 1,000 long texts from the RealDet dataset, truncated them to 300 tokens, and measured average detection cost with a batch size of 1. We observe that training-based methods such as Biscope and OpenAI-D were the fastest, requiring less than 0.1s per text, but these methods entail significant training overhead. Among training-free methods, classical methods like Likelihood, Logrank, and Entropy are faster, with inference times around 0.3s, but their detection accuracy did not meet our requirements. DNA-DetectLLM, Binoculars, and Fast-DetectGPT processed each sample in 0.8s, with DNA-DetectLLM achieving the better detection performance.
5 Conclusion
In this paper, we introduce DNA-DetectLLM, a novel zero-shot AI-generated text detection method via a DNA-inspired mutation-repair paradigm. Extensive experiments demonstrate that DNA-DetectLLM consistently achieves SOTA detection performance while exhibiting strong robustness across diverse scenarios. We hope our work offers new insights and perspectives for AI-generated text detection and plan to further explore the mutation-repair paradigm to enhance detection performance.
Acknowledgments
This work is supported by the Postdoctoral Fellowship Program of CPSF under Grant Number GZC20251076, and the National Natural Science Foundation of China (No.U2336202).
Appendix A Limitations
Due to memory constraints, we were unable to scale up the batch size for evaluating method efficiency under ample computational resources. As a result, this study does not fully explore performance differences in real-world scenarios involving large-scale, real-time monitoring of AI-generated texts. The reported efficiency results are based on relative comparisons under a uniform small-batch setting.
Appendix B Broader Impacts
The proposed DNA-DetectLLM contributes to the field of AI-generated text detection by improving accuracy and introducing a novel detection perspective to the research community. Its enhanced reliability may serve as a reference in socially relevant scenarios where auxiliary judgment is required. However, we emphasize that detection inherently carries a risk of implication or accusation. While our method demonstrates strong performance, we strongly oppose the use of its outputs as direct evidence in punitive or disciplinary contexts. Regardless of its accuracy, such applications could lead to serious consequences and misuse.
Appendix C Prompt Design for Main Data
Table 3: Examples of input prompts and corresponding outputs across different writing tasks, where the outputs are sampled from GPT-4-generated texts.
Writing Task Input Prompt Output
News Article Writing A police source told the BBC that an infiltrator from the Taliban had allowed militants into the police station in the regional capital of Lashkar Gah last night. Please continue. The militants, believed to be associated with the Taliban, reportedly raided the police station and engaged in a prolonged gun battle with the officers…
Story Generation Two kids entered the Rockmount Zoo. Please continue. As they excitedly crossed the threshold, their eyes widened at the sight of the vibrantly colored parrots squawking from the treetops…
Academic Writing Please write an abstract based on the following title: “Pure Exploration and Regret Minimization in Matching Bandits”. This paper delves into the field of pure exploration and regret minimization in the context of matching bandits problems - an important area in machine learning…
Table 3 presents the general-purpose prompts we designed for different writing tasks. Using these prompts, we generated 4,800 AI-generated texts—corresponding to human-written texts—across GPT-4 Turbo, Gemini-2.0 Flash, and Claude-3.7 Sonnet, which were used in our experiments. To ensure reproducibility, we explicitly report the generation parameters for each API call (note that Top-$k$ is not manually configurable):
  • GPT-4 Turbo: gpt-4-turbo-2024-04-09, Temperature = 1.0, Top-$p$ = 1.0.
  • Gemini 2.0 Flash: gemini-2.0-flash-001, Temperature = 1.0, Top-$p$ = 0.95.
  • Claude 3.7 Sonnet: claude-3-7-sonnet@20250219, Temperature = 1.0, Top-$p$ = 1.0.
Appendix D Main Experiments Supplement
All experiments are conducted on a single NVIDIA A100 GPU with 80GB of memory. Unless otherwise specified, default settings are used for temperature, top-k, and other generation parameters. No additional hyperparameter tuning is involved in this study. For training-based methods (Biscope and R-Detect), models are trained on 4,000 balanced samples from the HC3 dataset, and the best-performing checkpoints are selected based on validation performance on a separate 2,000-sample validation set. In the main experiments, all reported F1 scores are the maximum values obtained by selecting the optimal threshold based on the positive and negative scores of each method.
Table 4: Comparison of F1 score across public benchmark datasets.
Method M4 DetectRL Multi-LLM DetectRL Multi-Domain RealDet Avg.
OpenAI-D 68.01 70.55 67.89 70.41 69.22
Biscope 71.75 72.00 68.91 81.23 73.97
R-Detect 67.14 66.56 66.26 67.55 66.88
Entropy 75.39 70.10 55.04 60.33 65.22
Likelihood 66.82 66.62 66.60 66.87 66.73
LogRank 66.80 66.69 66.60 66.82 66.73
DetectGPT 54.53 42.41 30.79 66.59 48.58
Fast-DetectGPT 81.27 75.84 68.06 84.72 77.47
Binoculars 84.82 80.97 76.24 82.15 81.05
Lastde++ 82.74 69.17 61.72 84.59 74.56
DNA-DetectLLM 85.15 84.49 83.94 84.72 84.58
To ensure a fairer performance comparison, we select fixed thresholds for all methods based on scores computed on a separate clean dataset (e.g., DNA-DetectLLM: 0.6533, Binoculars: 0.9366, etc.). Subsequently, we recompute and report the F1 scores in Table 4 using these fixed thresholds. The clean dataset used for threshold selection consists of over 3,000 samples generated by GPT-4, Gemini, and Claude based on human-written texts sourced from XSum, WritingPrompt, and Arxiv.
Appendix E Additional Performance Comparisons
Table 5: Comparison of AUROC (%) across benchmark datasets.
Method XSum WritingPrompt Arxiv PubMedQA Avg.
Revise-Detect 39.73 65.54 95.31 66.86
GECScore 70.84 66.31 64.91 67.35
DNA-GPT 65.46 75.22 70.13 82.32 73.28
ImBD 88.07 93.06 91.06 92.59 91.20
GPTZero 99.01 98.54 94.42 88.48 95.11
DNA-DetectLLM 99.31 98.86 95.00 97.08 97.56
As shown in Table 5, we expanded our comparative experiments to include additional baselines—Revise-Detect [40], GECScore [34], DNA-GPT [38], IMBD [6] (recent but non–state-of-the-art methods), and GPTZero (a widely used commercial detector). We also incorporated results on the biomedical short-text dataset PubMedQA [19], which further demonstrate the strong and consistent performance of DNA-DetectLLM across diverse domains and detection settings.
Appendix F Robustness Experiments
In this study, we do not consider adversarial attacks on human-written texts, as evasion in such cases is generally inconsequential. Instead, we focus on adversarial scenarios involving AI-generated texts, introducing two common attack types: paraphrasing and token-level editing. For paraphrasing attacks, we employ DIPPER with hyperparameters set to a lexical diversity of 60 and a syntactic diversity of 60. This level of paraphrasing is sufficient to potentially bypass SOTA detectors. For editing attacks, we tokenize the input using the GPT-2 tokenizer and apply random insertions, deletions, and substitutions to 1% of the tokens. The inserted or substituted tokens are sampled uniformly from the tokenizer’s vocabulary.
Table 6: AUROC (%) for GPT-4 Turbo against various attacks.
Method Insertion Deletion Substitution Paraphrase
OpenAI-D 51.54 52.95 51.62 89.50
Biscope 60.75 62.66 58.59 41.16
R-Detect 61.71 61.27 61.08 57.49
Entropy 57.52 63.38 55.44 36.04
Likelihood 51.54 60.69 49.89 49.33
LogRank 52.19 60.69 50.45 50.48
DetectGPT 43.78 52.36 39.67 44.60
Fast-DetectGPT 86.16 92.05 85.89 96.70
Binoculars 87.55 93.32 87.28 97.23
Lastde++ 71.60 88.51 71.34 94.94
DNA-DetectLLM 93.37 96.28 93.06 98.02
Table 7: AUROC (%) for Gemini-2.0 Flash against various attacks.
Method Insertion Deletion Substitution Paraphrase
OpenAI-D 52.28 56.72 53.02 86.46
Biscope 95.48 95.79 95.15 52.61
R-Detect 56.40 56.37 56.84 52.97
Entropy 65.59 71.11 63.47 33.50
Likelihood 76.41 82.86 74.73 47.02
LogRank 76.79 82.93 74.98 47.65
DetectGPT 68.74 76.01 66.60 49.68
Fast-DetectGPT 95.39 96.57 95.23 95.27
Binoculars 97.41 98.06 97.35 96.78
Lastde++ 90.39 95.31 90.16 93.41
DNA-DetectLLM 97.72 98.16 97.69 97.80
Table 8: AUROC (%) for Claude-3.7 Sonnet against various attacks.
Method Insertion Deletion Substitution Paraphrase
OpenAI-D 57.99 60.50 57.80 83.62
Biscope 93.99 93.47 93.65 44.34
R-Detect 58.44 58.24 58.06 53.02
Entropy 73.32 78.29 70.87 35.93
Likelihood 64.11 73.01 61.93 42.80
LogRank 64.26 72.79 61.99 44.21
DetectGPT 60.39 68.81 57.54 45.37
Fast-DetectGPT 86.38 92.74 85.34 92.56
Binoculars 89.81 95.34 88.88 95.33
Lastde++ 73.47 89.85 72.54 90.43
DNA-DetectLLM 93.77 96.89 93.25 96.81
Table 9: F1 score (%) under paraphrasing and editing attacks
Method GPT-4 Turbo Gemini-2.0 Flash Claude-3.7 Sonnet Avg.
Insert Deletion Substitution Paraphrase Insert Deletion Substitution Paraphrase Insert Deletion Substitution Paraphrase
OpenAI-D 57.17 58.23 58.91 74.24 59.52 60.24 60.34 72.99 65.00 66.04 65.15 71.93 64.15
Biscope 54.31 55.19 50.76 27.09 86.46 86.34 86.30 42.42 86.09 85.80 85.80 32.73 64.94
R-Detect 66.46 66.46 66.46 66.50 66.16 66.16 66.16 66.16 66.63 66.63 66.63 66.63 66.42
Entropy 62.53 65.04 61.92 56.82 65.95 68.69 64.68 57.09 68.27 71.36 67.02 57.20 63.88
Likelihood 66.65 66.65 66.65 66.61 66.54 66.54 66.54 66.54 66.69 66.69 66.69 66.65 66.62
LogRank 66.61 66.61 66.61 66.61 66.57 66.57 66.57 66.50 66.65 66.65 66.65 66.65 66.60
DetectGPT 21.80 30.87 17.13 21.04 54.86 65.84 52.22 30.18 43.03 54.86 38.79 22.13 37.73
Fast-DetectGPT 77.10 84.28 76.24 89.38 87.90 89.04 87.85 88.03 77.00 85.22 75.81 84.80 83.55
Binoculars 81.33 87.16 81.15 91.50 92.44 93.08 92.28 91.65 83.15 88.71 82.36 88.78 87.80
Lastde++ 60.86 80.67 60.39 87.33 82.46 87.16 81.88 85.87 61.46 82.22 61.26 83.02 76.22
DNA-DetectLLM 86.74 90.36 86.01 93.09 94.22 94.94 93.91 93.63 87.28 91.58 86.81 91.06 90.80
Table 6, Table 7, and Table 8 report the detection performance of all methods under various adversarial attacks across different AI-generated texts. Notably, DNA-DetectLLM consistently achieves strong performance across all adversarial scenarios, demonstrating robustness to both editing and paraphrasing attacks. In contrast, training-free methods are significantly affected by token-level edits, while training-based methods are more vulnerable to paraphrasing-based attacks.
Table 9 reports the F1-score performances for the primary robustness experiments. Notably, DNA-DetectLLM consistently exhibits superior robustness compared to other baselines in practical scenarios.
Appendix G Ablation Study Supplement
Table 10: Ablation results across different repair strategies and datasets.
Setting XSum WP Arxiv M4 RealDet Avg. Time Cost (s)
Default 98.14 99.03 97.74 91.74 94.48 96.23 0.78
Low-to-High 97.08 98.49 96.71 92.42 94.71 95.88 14.11
High-to-Low 98.25 98.94 97.29 89.56 93.59 95.53 14.45
Sequential Repair 98.36 98.37 97.97 91.26 93.71 95.93 14.55
Table 10 compares the detection performance and inference time of DNA-DetectLLM under different repair orders. The results show that the default setting achieves superior performance compared to alternative strategies, while reducing computation time by nearly 20×, highlighting its practical efficiency.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The methodological pipeline and contributions of our approach, including the dataset construction, are clearly outlined in the abstract and introduction.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We discuss the limitations of this study in Appendix A.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: A complete derivation of the simplified repair score, including all assumptions, is provided in Section 3.6.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We fully disclose all information necessary to reproduce the main experimental results, including but not limited to the details provided in Sections 3,  4, and Appendix D.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in the supplemental material?
  23. Answer: [Yes]
  24. Justification: We have included the data and code in the supplementary material, along with detailed instructions to facilitate reproduction of the main experimental results.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We specify all the training and test details in Section 4.1.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: While we do not report error bars, we evaluate our method and all baselines using AUROC and F1 scores.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: Computational resources are described in detail in Appendix D.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We have thoroughly read the NeurIPS Code of Ethics and ensured compliance in all aspects of our research.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: We discuss the potential societal impacts of our work in Appendix B.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: Not involved in misusing.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We provide proper citations for all baseline methods and datasets used in this study.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: All code, datasets, and models are well-documented and openly available through access in the supplementary material.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: Our work does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: Our work does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [Yes]
  79. Justification: The LLMs used in this study are explicitly described in Sections 4.1 and  $4.$
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.